Consider a population consisting of 4, 7, 8, 13, 15 and 17. Suppose samples of size 3 are drawn from this population. Determine the total number of samples. *

1
Expert's answer
2022-04-14T03:15:08-0400

The number of possible samples which can be selected without replacement is

"\begin{pmatrix}
 N \\
 n
\end{pmatrix}=\cfrac{N! } {n! \cdot(N-n)! }=\\
=\begin{pmatrix}
 6\\
 3
\end{pmatrix}=\cfrac{6! } {3! \cdot3! }=\cfrac{4\cdot5\cdot6}{2\cdot3}=20 ."